Consider a collection of envelopes consisting of 3 red envelopes 3 blue envelopes​, 3 green envelopes an 2 yellow envelopes. If two envelopes are selected at​ random, with​ replacement, determine the probability that both are red envelopes.
Answer #1

P(both are red) = 3/11 * 3/11 = 9/121 = 0.0744